OpenStudy (chazz16):

Medal,Medal,Medal,Medal Will Be Awarded!!! Type the correct answer in each box. Use numerals instead of words. If necessary, use / for the fraction bar(s). The slope of diagonal PQ is , and its equation is . Graph Is In The Comments

OpenStudy (chazz16):

Graph

OpenStudy (retireed):

Slope = (Y2 - Y1) / (X2 - X1) numerator 2-2 + 0 denominator is 4 Slope = __________

OpenStudy (retireed):

Equation of the Line y =mx +b y = (0)x + 2 for the domain -?? <= x <= ???

OpenStudy (chazz16):

@retirEEd Can You Put It in This Format. The slope of diagonal PQ is_____________, and its equation is__________________ .

OpenStudy (retireed):

Figure out the slope and just fill in the blanks yourself. I already gave the equation you need to just figure out the domain.